OpenStudy (anonymous):

4. Solve the equation for x: 3x + 2 = 17 A. 5 B. 6 C. 7 D. 8

OpenStudy (anonymous):

3X= 15 X=15/3 X=5 A

OpenStudy (whpalmer4):

\[3x+2=17\]We want to isolate \(x\) all by itself on one side of the equals sign. To do that, we can add, subtract, multiply, divide, take square roots, etc. — we simply have to do the same thing to each side of the equals sign. What do you get if you subtract 2 from each side of this equation? \[3x+2-2 = 17-2\]\[3x=15\] That looks like progress. Now we divide both sides by 3, because that is the number that appears in front of \(x\) (its coefficient)\[\frac{3x}{3} = \frac{15}{3}\]\[x=5\]